# Neural Scaling Laws on Graphs

### v1

@@ Abstract
Deep graph models (e.g., graph neural networks and graph transformers) have become important techniques for leveraging knowledge across various types of graphs. Yet, the scaling properties of deep graph models have not been systematically investigated, casting doubt on the feasibility of achieving large graph models through enlarging the model and dataset sizes. In this work, we delve into neural scaling laws on graphs from both model and data perspectives. We first verify the validity of such laws on graphs, establishing formulations to describe the scaling behaviors. For model scaling, we investigate the phenomenon of scaling law collapse and identify overfitting as the potential reason. Moreover, we reveal that the model depth of deep graph models can impact the model scaling behaviors, which differ from observations in other domains such as CV and NLP. For data scaling, we suggest that the number of graphs can not effectively metric the graph data volume in scaling law since the sizes of different graphs are highly irregular. Instead, we reform the data scaling law with the number of edges as the metric to address the irregular graph sizes. We further demonstrate the reformed law offers a unified view of the data scaling behaviors for various fundamental graph tasks including node classification, link prediction, and graph classification. This work provides valuable insights into neural scaling laws on graphs, which can serve as an essential step toward large graph models.
Machine Learning, ICML
@@ 1 Introduction
Neural scaling laws (Rosenfeld et al., 2019; Kaplan et al., 2020; Abnar et al., 2021) quantitatively describe how model performance grows with the scale of training. There are two basic forms: data scaling law (how the performance changes with training dataset size) and model scaling law (how the performance changes with model size). Neural scaling laws guide the construction of large models from two perspectives: (1) They enable us to predict model performance on large-scale training by extrapolating from small-scale runs with lower computational budgets. For instance, the performance of GPT-4 on coding problems fits accurately with the prediction of the scaling law (Achiam et al., 2023). (2) They enable us to identify better model architectures through small-scale training, typically selecting those with the largest performance gains for a given computational overhead. Such laws have played an instrumental role in the recent successes of large models in natural language processing (NLP) and computer vision (CV) domains (Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023; Zhai et al., 2022; Dehghani et al., 2023).
@@
Despite these impacts on NLP and CV, no similar revolution has occurred in the graph machine learning domain (Zhang et al., 2023), given its unique difficulties. The investigation of neural scaling laws on graphs faces new challenges from both data and model perspectives. (1) From the data perspective: Graphs are irregular and diverse. Both the number of nodes and edges of a graph can vary dramatically. As an example, in the ogbg-ppa dataset of $158k$ graphs (Szklarczyk et al., 2019), the number of nodes in each graph ranges from $50$ to $300$, and the number of edges from $54$ to $18,096$, indicating significant variability across the dataset. The variability in size suggests that the current metric (graphs) insufficiently represents the data volume, hinting at the need for a new graph scaling law metric. (2) From the model perspective: Besides the model weights considered in the existing scaling law, deep graph models also need to consider the influence of the non-parametric aggregations, i.e., whether graph models with different numbers of aggregation layers could exhibit distinct scaling behaviors. This issue is not yet examined.
@@
In this work, we make a first attempt to address these unique challenges and investigate the model scaling law and the data scaling law on graphs. While the main focus of this work is on the graph classification tasks, we also show that our settings and results can be extrapolated to node classification and link prediction tasks. Experimental observations reveal neural scaling laws across different models and dataset sizes for graph learning tasks. Our contributions can be summarized as follows:
  • For the model scaling law: (i) We examine the model scaling law for graph learning, which reveals that performance scales with model size, but collapses when the model becomes excessively large, likely due to overfitting. This suggests graph learning models are likely to be data hungry. (ii) From the perspective of architecture, we find that deep graph models of different numbers of layers will exhibit different model scaling behaviors, contrasting with the conclusions in CV and NLP domains.
@@
  • For the data scaling law: (i) Regarding the size variability problem, we reveal that the number of edges is a better metric than the number of graphs when applying the data scaling law in graph classification tasks. To support this argument, we demonstrate that models trained on datasets with identical graph counts but different total numbers of edges exhibit distinctive scaling behaviors. (ii) We further extend the data scaling law to node classification and link prediction tasks, demonstrating the generality of using the number of edges as the data metric.
@@ 2 Related Work
Functional forms of neural scaling laws. Neural scaling laws give the quantitative relationship between the model performance and three main factors: the number of model parameters $N$, the size of the training dataset $D$, and the amount of computation $C$. The relationship is generally described as a power-law, which has widely been applied in the domains of image classification (Hestness et al., 2017; Henighan et al., 2020; Sharma & Kaplan, 2022; Zhai et al., 2022), generative language modeling (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022), and natural language translation (Gordon et al., 2021; Ghorbani et al., 2021; Bansal et al., 2022; Fernandes et al., 2023).
@@
Using the test error $\epsilon$ as a metric for model performance, Hestness et al. (2017) proposes the basic functional form of neural scaling laws as follows:
$\epsilon=aX^{-b}+\epsilon_{\infty}$ (1)
@@
The variable $X$ represents any of the parameters $N$, $D$, or $C$. The positive constants $a$ and $b$ govern the rate of error reduction as a function of $X$, while $\epsilon_{\infty}>0$ specifies the minimum achievable error. Equation 1 is appropriate for regression tasks but is ill-defined for classification tasks, since the error rate $\epsilon_{r}$ remains finite for models with zero training. To fix this problem, Zhai et al. (2022) proposes a modified function:
$\epsilon_{r}=a(X+c)^{-b}+\epsilon_{\infty}$ (2)
@@
where a positive shift coefficient $c$ is added to avoid having infinite $\epsilon_{r}$ when $X$ equals to zero. In this work, we will base our following investigations of graph classification tasks on this equation.
@@
The relationship between $N$, $D$ and overfitting. The negative exponent $-b$ in Equation 2 indicates that the performance gain will be diminishing if only one of the parameters $N$ and $D$ is increased. To break this bottleneck, the dataset size and the model size need to be enlarged together. Specifically, Kaplan et al. (2020) find that the model size and dataset size should increase together as $D\sim N^{\alpha}$ ($10<\alpha<$ 1) to attain optimal return for compute budgets. It is also observed that overfitting will likely occur if $N$ surpasses $D^{1/\alpha}$ significantly, which leads to negative performance gain.
@@
Graph learning and scaling. The graph is unique in representing non-Euclidean data with irregular and flexible structures, e.g., social networks (Yanardag & Vishwanathan, 2015), molecules (Gilmer et al., 2017), and protein networks (Szklarczyk et al., 2019). Despite the popularity of graph machine learning, neural scaling laws on graphs remain underexplored. Some works have looked into the neural scaling laws for molecule representation learning (Chen et al., 2023), but the exploration is limited to the data and model types. It is unclear how the conclusion can be generalized to other graphs. To the best of our knowledge, we are the first to investigate neural scaling laws for the general graph domain.
3 Preliminaries
@@
Figure 1: The model and data scaling behaviors of GIN on pcqm4mv2. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The lower the mean square error is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
(a) Model scaling on ogbg-ppa, $R^{2}=0.98$.
(b) Data scaling on ogbg-ppa, $R^{2}=0.98$.
Figure 2: The model and data scaling behaviors of GIN on ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The higher the accuracy is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
Orgnizations. The main focus of this work is on the two basic forms of neural scaling laws for the graph domain, i.e., the modeling scaling law and the data scaling law. To consolidate our conclusion, our investigations focus on the well-studied supervised graph learning while leaving the under-studied unsupervised learning tasks (Sun et al., 2022) as one future work. The following sections will be arranged as follows: Section 4 verifies the validity of the basic forms of neural scaling laws on the graphs. Section 5 investigates deeper into the model scaling law of graph models. We observe the model scaling collapse phenomenon when graph models are excessively large and identify overfitting as the potential reason. Moreover, we reveal that model depth can affect the model scaling behaviors, which is a unique phenomenon in graph domains. In Section 6, we propose the number of edges is a suitable data metric when applying the data scaling law on graphs. To support the argument, we exhibit that models would have distinctions in scaling behavior when trained on datasets with different numbers of edges. Furthermore, we present that using the number of edges as the data metric could provide a unified data scaling law for various graph learning tasks, e.g., node classification, link prediction, and graph property prediction.
[…]
$s=-a(X+c)^{b}+s_{\infty}$ (3)
where $s$ is the test score corresponding to specific metrics and $s_{\infty}=1-\epsilon_{\infty}$. Equation 3 will be used in the following sections to fit the scaling law on classification tasks. The details of datasets and experimental settings can be found in Appendix A.
4 The Validity of Basic Scaling Laws on Graphs
@@
We start our analysis by examining the basic forms of neural scaling laws in the graph domain. In other words, we aim to explore how the model performance changes with the dataset size and model size. Hence, we first evaluate the performance of the models across various sizes and training dataset scales. To simplify the experimental settings, we vary the model size by increasing the model width while fixing the model depth, and the effect of model depth will be further explored in Section 5. To obtain subsets with different sizes, we randomly subsample the original training set by different ratios. Then, we fit the scaling law Eq. 1 and 3 to the empirical performance data using the least square method. To measure the fitting quality, the coefficient of determination $R^{2}$ is adopted as a metric. In this subsection, we focus on GIN and leave the results of GCN in Appendix B. Our experiments are conducted on two large graph property prediction datasets: pcqm4mv2 and ogbg-ppa.
We conduct experiments for both model scaling and data scaling behaviors on the graph regression dataset pcqm4mv2. For regression tasks, the performance is measured by mean square error (MSE), hence we use Eq. 1 to fit the empirical MSEs. For model scaling, we fix the number of layers for the model to $6$ and vary the model width from $64$ to $1,024$. The experiments are repeated for different data fractions chosen from $\left[0.1,0.2,0.3,0.4,0.5,0.7,1\right]$. The results, shown in Figure 1(a), depict each model’s MSE as dots, with dashed lines representing fitted scaling law curves and colors indicating specific training data fractions. We make the following observations from Figure 1(a): (1) the MSE tends to decrease when increasing the model size and fixing the dataset size. (2) the MSE gradually decreases and approaches a lower bound $\epsilon$ asymptotically as the model becomes larger. The scaling curves corresponding to larger dataset sizes would have smaller $\epsilon$, i.e., higher upper bound of the model performance. The observations in Figure 1(b) exhibit similar patterns when scaling up data size while fixing the model size. Generally speaking, both model scaling curves in Figure 1(a) and data scaling curves in 1(b) fit well to empirical values with high $R^{2}$s, indicating that Eq. 1 could be leveraged to describe the neural scaling laws for graph regression tasks.
The scaling behaviors on the ogbg-ppa dataset are illustrated in Figure 2. The way of varying dataset sizes and model sizes is similar to that of pcqm4mv2. However, since the task on ogbg-ppa is graph classification, we use accuracy as the performance metric and apply Eq. 3 to fit the scaling behaviors. Observations from Figure 2 align with those from Figure 1, suggesting that Equation 3 can describe the scaling behavior for graph classification tasks.
Moreover, similar to Rosenfeld et al. (2019), the model scaling law and data scaling law formulations can be combined into a comprehensive form for regression tasks:
$\epsilon=a_{1}D^{b_{2}}+a_{2}N^{b_{2}}+\epsilon_{\infty}$ (4)
@@
The above observations suggest that the scaling behaviors for graph property prediction tasks follow neural scaling laws and can be correctly predicted by Eq. 1 and 3. Thus, it is possible to attain expected performance on graph tasks by enlarging the model and dataset size, exhibiting the potential of building large graph models.
(a) Model scaling on ogbg-molpcba
(b) Training losses of models
(c) Validation losses of models
Figure 4: (a) The empirical performance of larger models falls below the prediction of model scaling law. (b)(c) The training losses and validation losses of models with varying sizes. The color corresponds to the number of model parameters. The model with larger sizes suffers more serious overfitting
5 Graph-specific Phenomena for Model Scaling Laws
In this section, we will dive deeper into specific phenomena of model scaling on graphs. In Section 5.1, we observe the collapse of model scaling in certain situations and discuss the potential reason. In Section 5.2, we identify that models with different depths could exhibit distinct model scaling behaviors. These observations enable us to have a better understanding of the uniqueness of model scaling law on graphs, providing important instructions for developing large graph models.
5.1 Model Scaling Collapse
The model scaling collapse poses difficulty in building large graph models following the model scaling law. Thus, it is essential to explore why the collapse happens and how to mitigate it. In this work, we conjecture that one of the potential reasons for the scaling law collapse is overfitting. To dive deeper into the problem, Figure 4(b) and 4(c) present the training loss and validation loss curves, respectively. All the training losses in Figure 4(b) keep dropping in the training process. The larger the model is, the smaller the corresponding training loss will be. Nonetheless, the validation losses of models with more than $10^{7}$ parameters start increasing after several epochs as shown in Figure 4(c), which indicates overfitting. Moreover, the models whose performance falls behind the scaling law predictions suffer more serious overfitting. For instance, models of 7.4M, 13M, and 20M parameters overfit the most in Figure 4(c), and they are the three models on which model scaling law fails in Figure 4(a). Based on these observations, overfitting can be one potential reason for the model scaling collapse. For more results between the model size and overfitting, please refer to Appendix C.
(a) Model scaling of GCN with different layers.
(b) Model scaling of GIN with different layers.
Figure 6: Model scaling behaviors of GNNs with varying model depths. The color corresponds to the model layer. Both GIN and GCN exhibit distinct scaling behaviors with different model depths.
(a) Model scaling of SAT with different layers.
(b) Model scaling of GraphGPS with different layers.
Figure 7: Model scaling behaviors of graph transformers with varying model depths. The color corresponds to the model layer. Both SAT and GraphGPS exhibit distinct scaling behaviors with different model depths.
To eliminate the effect of model scaling collapse, it is crucial to overcome the overfitting. Kaplan et al. (2020) suggest that overfitting would happen if the model size is too large for the training set. To investigate whether more training data can mitigate overfitting, we train a model of fixed number parameters on different fractions of the training set and illustrate the validation loss curves in Figure 5. The model trained on $10$ percent of the original ogbg-molpcba dataset suffers serious overfitting as the validation loss curve suggests. However, the overfitting is significantly mitigated when the data fraction is raised to $50$ percent, which indicates the effectiveness of enlarging the training set. Another evidence is that we have not observed overfitting or model scaling collapse in models of similar sizes ($10^{6}$ parameters) trained on larger datasets, i.e., pcqm4mv2 and ogbg-ppa in Section 4.
Observation 2. The model scaling collapses in the graph domain often happen due to the overfitting.
5.2 The Effect of Model Depth

### v3

# Towards Neural Scaling Laws on Graphs

@@ Abstract
Deep graph models (e.g., graph neural networks and graph transformers) have become important techniques for leveraging knowledge across various types of graphs. Yet, the neural scaling laws on graphs, i.e., how the performance of deep graph models changes with model and dataset sizes, have not been systematically investigated, casting doubts on the feasibility of achieving large graph models. To fill this gap, we benchmark many graph datasets from different tasks and make an attempt to establish the neural scaling laws on graphs from both model and data perspectives. The model size we investigated is up to 100 million parameters, and the dataset size investigated is up to 50 million samples. We first verify the validity of such laws on graphs, establishing proper formulations to describe the scaling behaviors. For model scaling, we identify that despite the parameter numbers, the model depth also plays an important role in affecting the model scaling behaviors, which differs from observations in other domains such as CV and NLP. For data scaling, we suggest that the number of graphs can not effectively measure the graph data volume in scaling law since the sizes of different graphs are highly irregular. Instead, we reform the data scaling law with the number of nodes or edges as the metric to address the irregular graph sizes. We further demonstrate that the reformed law offers a unified view of the data scaling behaviors for various fundamental graph tasks including node classification, link prediction, and graph classification. This work provides valuable insights into neural scaling laws on graphs, which can serve as an important tool for collecting new graph data and developing large graph models.
@@ 1 Introduction
The neural scaling law [1, 2, 3] serves as an important tool in the development of large models. It quantitatively describes how model performance grows with the scale of training, which have two basic forms: data scaling law (how the performance changes with training dataset size) and model scaling law (how the performance changes with model size). Neural scaling laws guide the construction of large models from the following perspectives: (1) They enable us to predict model performance on large-scale training by extrapolating from small-scale runs with lower computational budgets. For instance, the performance of GPT-4 on coding problems fits accurately with the prediction of the scaling law [4]. (2) They enable us to identify better model architectures through small-scale training, typically selecting those with the largest performance gains for a given computational overhead. (3) They serves as the guidance for resource allocation, i.e., they could tell us how to distribute the cost of data collection and computation for the given model to reach the expected performance. Such laws have played an instrumental role in the recent successes of large models in natural language processing (NLP) and computer vision (CV) domains [5, 6, 7, 8, 9].
Figure 1: The density distribution of the number of nodes and edges in ogbg-ppa dataset.
@@
Despite the success of foundation models in NLP and CV, there is no existing effort investigating the neural scaling law in the graph machine learning domain. As an initial exploration, we found that such an investigation would meet unique challenges given the irregularity and diversity of graphs. Specifically, both the number of nodes and edges of a graph can vary dramatically, different from the fixed-structure figures or texts in CV and NLP. As an example shown in Figure 1, in the ogbg-ppa dataset of $158k$ graphs [10], the number of nodes in each graph ranges from $25$ to $325$, and the number of edges from $54$ to $18,096$, indicating significant variability across the dataset. The variability in size suggests that the current metric (the number of graph samples) insufficiently represents the data volume, hinting at the need for more proper data metrics of the graph scaling laws. Moreover, the irregularity of graphs requires the deep graph models to introduce the message aggregation mechanisms. Hence, besides the model weights considered in the existing scaling law, deep graph models also need to consider the influence of the non-parametric aggregations, i.e., whether graph models with different numbers of aggregation layers could exhibit distinct scaling behaviors.
@@
In this work, we make a first attempt to address these unique challenges and develop the neural scaling laws on graphs to become a tool for building large graph models. Our investigation comes across a large range of domains and tasks in graph representation learning. Experimental observations reveal neural scaling laws across different models and dataset sizes for graph learning tasks.
Our contributions can be summarized as follows:
  • We start our investigation with verifying the general forms of graph scaling laws. Specifically, we develop math formulas for the relation between model/dataset size and the model performance. Then we verify their validity with empirical studies across different datasets and models.
  • We find that deep graph models with different numbers of layers will exhibit different model scaling behaviors, contrasting with the conclusions in CV and NLP domains. We then look deeper into the phenomenon and propose that the non-parametric aggregation operation could be one of the possible reasons.
@@
  • Regarding the size variability problem, we reveal that the number of nodes or edges is a better metric than the number of graphs when applying the data scaling law in graph classification tasks. To support this argument, we demonstrate that models trained on datasets with identical graph counts but different total numbers of edges exhibit distinctive scaling behaviors.
  • We further extend the data scaling law to node classification and link prediction tasks, demonstrating the generality of using the number of nodes or edges as the data metric.
Orgnizations. The main focus of this work is on the two basic forms of neural scaling laws for the graph domain, i.e., the modeling scaling law and the data scaling law. We tend to address the challenges caused by the irregularity of graph data when establishing the scaling laws. Our investigations focus on the well-studied supervised graph learning while leaving the under-studied unsupervised learning tasks [11] as one future work. The following sections will be arranged as follows: In Section 2, we briefly introduce the neural scaling laws and the related works. In Section 3, we introduce the basic settings of our experiments. In Section 4, we verify the validity of the basic forms of neural scaling laws on the graphs. In Section 5, we reveal that model depth can affect the model scaling behaviors, which is a unique phenomenon in graph domains. In Section 6, we propose the number of edges is a suitable data metric when applying the data scaling law on graphs. To support the argument, we exhibit that models would have distinctions in scaling behavior when trained on datasets with different numbers of edges. Furthermore, we present that using the number of edges as the data metric could provide a unified data scaling law for various graph learning tasks, e.g., node classification, link prediction, and graph property prediction.
@@ 2 Related Work
Functional forms of neural scaling laws. Neural scaling laws give the quantitative relationship between the model performance and three main factors: the number of model parameters $\mathbf{N}$, the size of the training dataset $\mathbf{D}$, and the amount of computation $\mathbf{C}$. The relationship is generally described as a power-law, which has widely been applied in the domains of image classification [12, 13, 14, 8], generative language modeling [2, 13, 15], and natural language translation [16, 17, 18, 19].
@@
Using the test error $\mathbf{\epsilon}$ as a metric for model performance, Hestness et al [12] propose the basic functional form of neural scaling laws as follows:
$\mathbf{\epsilon=aX^{-b}+\epsilon_{\infty}}$ (1)
@@
The variable $\mathbf{X}$ represents any of the parameters $\mathbf{N}$, $\mathbf{D}$, or $\mathbf{C}$. The positive constants $\mathbf{a}$ and $\mathbf{b}$ govern the rate of error reduction as a function of $\mathbf{X}$, while $\mathbf{\epsilon_{\infty}}>0$ specifies the minimum achievable error. Equation 1 is appropriate for regression tasks but is ill-defined for classification tasks, since the error rate $\mathbf{\epsilon_{r}}$ remains finite for models with zero training. To fix this problem, Zhai et al [8] proposes a modified function:
$\mathbf{\epsilon_{r}=a(X+c)^{-b}+\epsilon_{\infty}}$ (2)
@@
where a positive shift coefficient $\mathbf{c}$ is added to avoid having infinite $\mathbf{\epsilon_{r}}$ when $\mathbf{X}$ equals to zero. In this work, we will base our following investigations of graph classification tasks on this equation.
@@
The relationship between $\mathbf{N}$, $\mathbf{D}$ and overfitting. The negative exponent $-\mathbf{b}$ in Equation 2 indicates that the performance gain will be diminishing if only one of the parameters $\mathbf{N}$ and $\mathbf{D}$ is increased. To break this bottleneck, the dataset size and the model size need to be enlarged together. Specifically, Kaplan et al [2] find that the model size and dataset size should increase together as $\mathbf{D\sim N^{\alpha}}$ ($10<\mathbf{\alpha}<$ 1) to attain optimal return for compute budgets. It is also observed that overfitting will likely occur if $\mathbf{N}$ surpasses $\mathbf{D^{1/\alpha}}$ significantly, which leads to negative performance gain.
@@
Graph learning and scaling. The graph is capable of representing non-Euclidean data with irregular and flexible structures, e.g., social networks [20], molecules [21], and protein networks [10]. Despite the popularity of graph machine learning, there are seldom large graph models. Notably, most deep GNNs remain small parameter size but more aggregation layers. Moreover, the neural scaling laws on graphs remain underexplored. Some works have looked into the neural scaling laws for molecule representation learning [22, 23], but the exploration is limited to the data and model types. It is unclear how the conclusion can be generalized to other graphs. To the best of our knowledge, we are the first to investigate neural scaling laws for the general graph domain.
3 Preliminaries and Experiment Setups
To ensure the applicability of neural scaling laws in the graph domain, our experiments are comprehensive to cover different types of models and graphs from various domains, e.g., molecule, biology, and academic. We investigate two major types of deep graph learning models, i.e., graph neural networks (GNNs) and graph transformers. We explore GCN [24] and GIN [25] for GNNs, and SAT [26], and GraphGPS [27] for graph transformers. Our datasets [28, 29, 30] cover the the most common graph domains including social networks (reddit-multi-12k, ogbl-collab), molecules (ogbg-molhiv,ogbg-molpcba,pcqm4mv2), protein networks (ogbg-ppa), and citation networks (ogbn-arxiv). Also, the chosen datasets are relatively large-scale ones among the academic datasets, ensuring valid observation of scaling behaviors. Moreover, our experiments are conducted on three common graph tasks, i.e. graph property prediction, node classification, and link prediction, though graph property prediction is the focus. Notably, accuracy rather than error rate is adopted in graph classification tasks as the default metric. To ensure a consistent observation, we switch the error rate $\epsilon_{r}$ in Equation 2 to accuracy via applying the relation $\text{Accuracy}=1-\text{ErrorRate}$ and obtain the following equation:
$\mathbf{s=-a(X+c)^{-b}+s_{\infty}}$ (3)
where $\mathbf{s}$ is the test score corresponding to specific metrics and $\mathbf{s_{\infty}=1-\epsilon_{\infty}}$. Equation 3 will be used in the following sections to fit the scaling law on classification tasks. The details of datasets and experimental settings can be found in Appendix A.
4 Validity of Basic Scaling Laws on Graphs
We start our analysis by examining the basic forms of neural scaling laws in the graph domain. In other words, we aim to explore whether the model performance has a power-law relationship with the dataset size and model size. Hence, we first evaluate the performance of the models across various sizes and training set sizes. Since the impact model depth is unclear yet, we vary the model size by
increasing the model width while fixing the model depth, and the effect of model depth will be further explored in Section 5. To obtain subsets with different sizes, we randomly subsample the original training set by different sampling ratios. Then, we fit the scaling law Equation 1 and 3 to the empirical results on different model scales using the least square method. To measure the fitting quality, the coefficient of determination $R^{2}$ is adopted as a metric.
In this subsection, we focus on GIN and leave the results with similar observations on graph transformer in Section 5 on GCN in Appendix B. Our experiments are conducted on two large graph property prediction datasets: pcqm4mv2 and ogbg-ppa.
@@ 4 Validity of Basic Scaling Laws on Graphs
(c) Model scaling on ogbg-ppa, $R^{2}=0.98$.
(d) Data scaling on ogbg-ppa, $R^{2}=0.98$.
Figure 2: The model and data scaling behaviors of GIN on both pcqm4mv2 and ogbg-ppa. For model scaling, the color corresponds to the training data fraction; for data scaling, the color corresponds to the number of model parameters. The lower the mean square error is, the better the model performance is. The high $R^{2}$ values indicate that the scaling law curves fit well with the empirical values.
We conduct experiments for both model scaling and data scaling behaviors on the graph regression dataset pcqm4mv2. For regression tasks, the performance is measured by mean square error (MSE), hence we use Equation 1 to fit the empirical MSEs. For model scaling, we fix the number of layers for the model to $6$ and vary the model width from $64$ to $1,024$. The experiments are repeated for different data fractions chosen from $\left[0.1,0.2,0.3,0.4,0.5,0.7,1\right]$. The results, shown in Figure 2(a), depict each model’s MSE as dots, with dashed lines representing fitted scaling law curves and colors indicating specific training data fractions. We make the following observations from Figure 2(a): (1) The MSE tends to decrease when increasing the model scale and fixing the dataset scale. (2) The MSE gradually decreases and approaches a lower bound $\epsilon$ asymptotically as the model becomes larger. The scaling curves corresponding to larger dataset sizes would have smaller $\epsilon$, i.e., higher upper bound of the model performance. The observations in Figure 2(b) exhibit similar patterns when scaling up data size while fixing the model size. Generally speaking, both model scaling curves in Figure 2(a) and data scaling curves in Figure 2(b) fit well to empirical values with high $R^{2}$s, indicating that Equation 1 could be leveraged to describe the neural scaling laws for graph regression tasks.
@@
More scaling behaviors on the ogbg-ppa dataset are illustrated in Figure 2(c) and 2(d). The way of varying dataset sizes and model sizes is similar to that of pcqm4mv2. However, since the task on ogbg-ppa is graph classification, we use accuracy as the performance metric and apply Equation 3 to fit the scaling behaviors. Observations from Figure 2(c) and 2(d) align with those from Figure 2(a) and 2(b), suggesting that Equation 3 can describe the scaling behavior for graph classification tasks.
On the basis of the validity of data scaling law and model scaling law individually, we further show that the two formulations can be combined into a comprehensive form, similar to [1]. For regression, the form is:
$\mathbf{\epsilon=a_{1}D^{-b_{2}}+a_{2}N^{-b_{2}}+\epsilon_{\infty}}$ (4)
@@
The above observations suggest that the scaling behaviors for graph property prediction tasks follow neural scaling laws and can be correctly predicted by Equation 1 and 3. Thus, it is possible to attain expected performance on graph tasks by enlarging the model and dataset size, exhibiting the potential of building large graph models. Despite the general success, there are still some unique chanllenges for scaling laws in the graph domain. We will discuss the challenges for the model scaling in Section 5, and those for the data scaling in Section 6.
5 The Effect of Model Depth on Graph Scaling Laws